OpenStudy (anonymous):

Write an equation of the line. Slope 0; through (-2,-1) The equation of the line is

OpenStudy (anonymous):

\[y-y1=m(x-x1)\] apply this formula in all; these types of questions @MarthaHearst

OpenStudy (anonymous):

Then what @rosebird

OpenStudy (anonymous):

y + 1 = 0 (x + 2) y = -1

OpenStudy (anonymous):

yeh y=-1

OpenStudy (anonymous):

@MarthaHearst